Identify each compound in the following questions and make assignments in the 1H NMR

g) Compound 7 molecular formula of C10Ha5N and 1H NMR given. 6H 4H 2Н Зн 10 9 8 7 6 3 т 2 1 0 5 ppm
Answer #1

H NMR Cotign D.BE - CA- CH-~> +) - 10-(5) +1 = 10-14 +1 = 1)-7=4 Benzene ring is present f To-7.2 pm; sultiplet, ah, aromat